# Hang Xu3,   Hanhui Li1,   Mrinmaya Sachan2,5,   Xiaodan Liang1

### v7

@@ Abstract
We present SeePhys, a large-scale multimodal benchmark for LLM reasoning grounded in physics questions spanning across middle school to PhD qualifying exams. The benchmark covers 7 fundamental physics domains, incorporating 21 categories of highly heterogeneous diagrams. In contrast to prior works where visual elements mainly serve supplementary purposes, our benchmark features a substantial proportion of vision-essential problems (75%) that mandate visual information extraction for problem-solving. Through extensive evaluation, we observe that even the best-performing reasoning models (e.g., Gemini-2.5-pro and OpenAI-o4-mini) achieve sub-60% accuracy on our benchmark. These results reveal fundamental challenges in current large language models’ visual understanding capabilities, particularly in: (i) establishing meaningful connections between diagram interpretation and physics reasoning, and (ii) overcoming their persistent reliance on textual cues as cognitive shortcuts.
@@ 1 Introduction
While mathematical reasoning has been a cornerstone for evaluating the reasoning capability of large language models (LLMs) and multimodal large language models (MLLMs) GSM8K ; MATH ; Omni-MATH ; CombiBench ; MathVista ; chen2021geoqa ; lu2021iconqa ; chen2022unigeo ; MathVerse ; MATH-Vision ; AtomThink , the realm of natural science, especially physics, remains understudied as a much larger, even more diverse testbed for complex scientific reasoning. Physics reasoning inherently binds text explanations to real-world visual contexts, exposing critical gaps in their ability to emulate human-like world modeling PHYBench ; UGPhysics ; OlympiadBench ; PHYSICS .
@@
Frontier models have begun to demonstrate abstract perception of physical scenarios and logical reasoning capabilities Deepseek-r1 ; o1 ; o3_o4-mini ; Gemini-pro . Due to the intrinsic diversity of physics diagrams and the fact that they inherently reflect the law of physics in our world, developing a comprehensive benchmark for evaluating physics reasoning abilities and cross-modal understanding is crucial for enhancing LLMs.
@@
Early research primarily focused on assessing physical commonsense Physion and basic scientific knowledge ScienceQA , which was later gradually extended to competition-level and university-level physics problems GPQA ; PHYSICS . Due to the broad knowledge scope and high annotation difficulty inherent, some studies only considered test samples from limited knowledge levels PhysReason ; OlympiadBench ; MMMU . Furthermore, other works targeted the evaluation of language models and did not incorporate visual information PHYBench ; PHYSICS ; UGPhysics . Notably, compared to the mathematics domain, there has been limited exploration of models’ perceptual differences in processing physics diagrams, despite their richer topological structures.
@@
Figure 1: Overview of SeePhys. It encompasses 7 core physics domains and 21 diagram types, spanning the full knowledge spectrum from middle school to PhD qualifying exams levels.
@@
With these challenges in mind, we introduce SeePhys, to measure LLMs’ capability to visually understand the law of physics. It is a fully multimodal benchmark for evaluating physics reasoning across middle school to PhD qualifying exam levels. SeePhys comprises 2,000 rigorously validated questions collected from open-source textbooks, exercises, examinations, and competitions. These questions span 7 major fields of both classical and modern physics. To assess the extent to which different models rely on visual information for reasoning, we curate two subsets with different degrees of visual information enrichment and additionally compile supplementary copies of 2,000 purely visual instances where all problem statements in texts are presented in picture form. Through meticulous selection of 21 diagram types by domain experts, each problem challenges frontier MLLMs to integrate domain knowledge with visual understanding of physics diagrams (e.g., Feynman diagrams for particle interactions and Circuit diagrams for Electromagnetism). With SeePhys, we conduct extensive experiments to evaluate 28 leading LLMs and MLLMs such as o4-mini o3_o4-mini and Gemini-2.5-Pro Gemini-pro . The results reveal that even with extensive chain-of-thought, none of the current models could surpass 55% accuracy. Our analysis reveals that even non-essential diagrams can enhance physics reasoning performance when presented to MLLMs. We have also raised a challenge [^1] https://www.codabench.org/competitions/7925/ in the 2nd AI for Math Workshop at ICML 2025222https://sites.google.com/view/ai4mathworkshopicml2025/.
@@ 2 Related Work
Math reasoning benchmarks. Mathematical reasoning has emerged as a central testbed for evaluating LLMs. Early benchmarks like GSM8K GSM8K established the foundation for multi-step textual reasoning through elementary problems, while MATH MATH advanced the field with competition-level tasks (e.g., AMC/AIME), exposing critical limitations of early models. As these benchmarks achieved saturation, the community shifted toward higher complexity, e.g., introducing Olympiad-level challenges requiring formal theorem proving and combinatorial reasoning Omni-MATH ; CombiBench . Concurrently, the rise of multimodal reasoning spurred benchmarks such as MathVista MathVista and MATH-V MATH-Vision to integrate visual comprehension (e.g., diagrams, plots) with compositional reasoning. However, MathVerse MathVerse has found that MLLMs tend to rely on the reasoning capabilities of language models when performing mathematical tasks. In contrast, scientific diagrams, which are abstractions derived from real-world scenarios, with their complex visual features, may provide a more effective testbed for benchmarking models’ multimodal capabilities.
Physics Reasoning benchmarks. Contemporary physics benchmarks target broad scope of domains, with frontier datasets now encompassing: (1) PhD-candidate-level theoretical problems, and (2) Olympiad-style multi-step reasoning challenges requiring specialized physics intuition. Text-based physics benchmarks like PHYBench PHYBench , TPBench TPBench , and UGPhysics UGPhysics provide challenging problems that test advanced reasoning skills, yet fundamentally lack the visual components necessary to assess diagram interpretation abilities. Multimodal physics benchmarks such as PhysReason PhysReason , OlympiadBench OlympiadBench , and PHYSICS PHYSICS emphasize visual reasoning challenges without analysis regarding the extent of visual components’ impact. Moreover, constrained by the high annotation costs stemming from the extensive domain knowledge required and the scarcity of qualified visual materials, these datasets lack comprehensive coverage across knowledge hierarchies and detailed annotations of diagram types. To address these limitations, this paper contributes a dataset with multimodal and full-spectrum physics problems.
@@ 3.1 Data Collection Principles
UGPhysics UGPhysics 0 11,040
PHYSICS PHYSICS * 298 1,297
PHYBench PHYBench 0 500
PhysReason PhysReason * 972 1,200
TPBench TPBench 0 57
@@
ScienceQA ScienceQA 2,797 3,215
OlympiadBench OlympiadBench 1,958 2,334
SciBench SciBench * 177 291
SciEval SciEval 0 1,657
MMMU MMMU 444 443
MMMU-Pro MMMU-Pro 65 60
GPQA GPQA 0 227
ARB ARB 31 129
HLE HLE 28 230
@@
Wide knowledge spectrum. To provide a comprehensive evaluation of the model’s physics knowledge comprehension and modeling capabilities, questions are systematically selected across the following eight progressive knowledge levels: middle school, high school, beginner Olympiad problems, advanced Olympiad problems, undergraduate, senior undergraduate, master’s, and PhD qualifying exams. Domain experts in physics are invited to selectively curate the most representative problems at each knowledge tier.
@@ 3.2 Annotation
We then utilize Mathpix [^3] https://mathpix.com/ to perform OCR parsing on the collected documents and convert them into markdown text. Afterwards, we construct the (question, reasoning, answer) triples for each question. GPT-4.1 GPT-4.1 is employed to process redundant line breaks, string omissions, and LaTeX syntax errors. Finally, the trained annotators use a LaTeX parser to conduct manual verification on the triples.
@@
Data leakage prevention. To minimize the risk of data leakage, we eliminate samples with inconsistent responses by toggling the search function of GPT-4o GPT-4o on and off. Inspired by Rein et al. GPQA , we subsequently conduct a manual search for the remaining questions with correct responses.
Multimodal enhancement. To further eliminate the influence of textual modality information, we first use o4-mini o3_o4-mini to add a caption field to each example, containing a detailed description of geometric features and numerical information. We then augment the original 2,000 questions to obtain a purely visual version. Specifically, we render each question along with its corresponding diagrams into a single image up to 4,096×4,096 pixels. During rendering, we introduce variations in font types and sizes while ensuring readability based on the diagram’s dimensions. The introduction of purely visual QA not only enhances the authenticity of evaluation but also advances the model’s cross-modal understanding capabilities—mirroring human cognition by extracting key features from pixels, associating abstract concepts, and ultimately achieving problem-solving accuracy comparable to that under text-image QA conditions. Appendix A demonstrates the cases.
@@ 4.1 Evaluation Protocol
Composite judgment strategy. Recent advancements in the reasoning capabilities of LLMs enable them to discern key information within complex responses, thereby mitigating the inaccuracies often introduced by applying template matching to open-ended questions. Therefore, we further develop a composite judgment strategy based on a combination of LLM and template matching. As a first step, the model generates a response to the given input question and significant figures. Subsequently, the final answer is extracted through template matching and LLM-based processing. During the scoring process, SymPy [^4] https://www.sympy.org is first utilized to perform an initial screening for straightforward answers. Responses that do not pass the screening are then compared with the ground-truth using LLM. We apply accuracy as the metric for this deterministic evaluation. In the experiments in this paper, we use DeepSeek-V3 Deepseek-v3 as the extraction and judge model.
@@ 4.2 Evaluation Models
9 Large language models: DeepSeek-R1 Deepseek-r1 , DeepSeek-V3 Deepseek-v3 , Qwen3-235B-A22B Qwen3 , Qwen2.5-72B-Instruct Qwen2.5 , QwQ-32B QwQ-32B , R1-Distilled-Llama-70B Deepseek-r1 , Llama-4-Scout-17B Llama4 , Gemma3-27B Gemma3 , Llama-3.1-8B Llama-3.1 . We evaluate them with the TC setting.
19 Multimodal large language models: OpenAI o4-mini o3_o4-mini , OpenAI o3-mini o3-mini , OpenAI o1 o1 , Gemini-2.5-Pro Gemini-pro , Claude 3.7 Sonnet Claude-3.7-Sonnet , Doubao-1.5-pro Doubao-1.5-pro , GPT-4.1 GPT-4.1 , GPT-4o GPT-4o , QvQ-72B-preview QVQ-72b-preview , Qwen-VL series Qwen2.5-VL ; Qwen2-VL , Llama-3.2-Vision series Llama-3.2-11B-Vision , LLaVA-NeXT-7B LLaVA-NeXT-7B , Phi-4-multimodal Phi-4-multimodal , InternVL2.5-8B InternVL2.5 , LLaVA-OneVision-7B Llava-onevision . All these MLLMs are benchmarked with the TV/TC/TO/VO settings.
@@ 4.3 Performance across Differential Knowledge Levels
DeepSeek-R1 Deepseek-r1 54.9 46.9 47.7 31.9 49.9 34.2 49.0 41.2 42.2
DeepSeek-V3 Deepseek-v3 53.9 42.6 36.4 22.8 45.4 29.7 35.9 37.5 36.0
R1-Distilled-Llama-70B Deepseek-r1 48.0 41.4 34.6 14.2 31.5 16.0 28.9 25.9 26.9
Qwen3-235B-A22B Qwen3 47.1 33.7 31.8 20.4 41.2 25.1 31.7 30.7 31.1
Qwen2.5-72B-Inst Qwen2.5 41.2 40.2 25.2 8.2 26.8 12.8 18.6 17.8 21.1
QwQ-32B QwQ-32B 47.1 42.2 44.9 15.5 40.0 20.1 32.4 24.0 29.7
Llama-4-Scout-17B Llama4 48.0 36.5 31.8 11.3 28.5 14.2 28.3 26.1 24.8
Gemma3-27B Gemma3 21.6 36.5 30.8 5.1 23.1 9.1 15.2 11.9 16.9
Llama-3.1-8B Llama-3.1 26.5 15.7 17.8 3.9 7.6 3.7 10.3 8.4 9.2
@@
o4-mini o3_o4-mini 66.7 61.8 56.1 41.8 53.8 45.7 51.0 53.4 51.9
o3-mini o3-mini 47.1 46.2 39.3 28.3 47.0 36.1 48.3 42.3 40.3
o1 o1 60.8 56.6 50.5 32.5 54.4 40.6 52.4 40.4 45.6
Gemini-2.5-Pro Gemini-pro 69.6 66.7 64.5 46.7 64.2 50.2 53.8 44.2 54.9
Claude-3.7-Sonnet Claude-3.7-Sonnet 52.9 51.8 43.0 16.7 41.4 26.5 33.8 32.4 34.6
Doubao-1.5-pro Doubao-1.5-pro 70.6 58.2 49.5 29.2 56.6 34.7 40.7 37.5 43.9
GPT-4.1 GPT-4.1 51.0 52.6 41.1 17.0 39.7 31.1 42.1 35.6 35.3
GPT-4o GPT-4o 37.3 39.0 34.6 7.5 23.4 15.5 24.1 21.8 21.9
QVQ-72b-preview QVQ-72b-preview 38.2 36.5 30.8 11.3 25.9 14.2 26.2 20.2 22.5
Qwen2.5-VL-72B-Inst Qwen2.5-VL 61.8 42.2 29.0 10.4 29.9 14.6 18.6 19.4 24.2
Qwen2.5-VL-7B-Inst Qwen2.5-VL 39.2 25.3 21.5 4.2 8.7 5.9 10.3 7.3 11.6
Qwen2.5-VL-3B-Inst Qwen2.5-VL 30.4 21.3 13.1 2.9 10.4 7.3 6.2 6.2 9.8
Qwen2-VL-7B-Inst Qwen2-VL 24.5 17.3 14.0 4.4 8.5 4.6 10.3 7.0 9.2
Llama-3.2-90B-Vision Llama-3.2-11B-Vision 21.6 25.7 22.4 3.9 9.3 10.0 12.4 8.9 11.7
Llama3.2-11B-Vision Llama-3.2-11B-Vision 23.5 18.5 14.0 4.2 5.4 3.7 4.8 7.5 8.3
LLaVA-NeXT-7B LLaVA-NeXT-7B 14.5 12.7 11.2 5.5 13.2 8.2 11.0 9.4 8.7
Phi-4-multimodal Phi-4-multimodal 20.6 12.4 12.1 4.4 7.0 5.0 8.3 4.9 7.6
InternVL2.5-8B InternVL2.5 17.6 12.4 9.3 2.9 5.6 3.2 4.1 5.1 6.2
LLaVA-OneVision-7B Llava-onevision 20.6 10.8 12.1 2.7 5.4 2.3 6.2 5.4 6.1
@@
Diminishing returns of knowledge injection. Table 2 further illustrates performance disparities across models at varying knowledge levels. Contrary to expectations, the difficulty progression for models does not strictly follow the knowledge level (e.g., senior undergraduate and advanced Olympiad questions exhibit lower accuracy than PhD qualifying exams). This discrepancy suggests that current models primarily rely on knowledge memorization rather than truly learning the derivation patterns of scientific laws. Furthermore, by examining results from middle school to PhD, we observe that weaker models demonstrate a substantially steeper performance reduction ratio (e.g., LLaVA-OneVision-7B, -74.5%) compared to stronger ones (e.g., o4-mini, -19.9%). Not only indicates that cutting-edge models fail to grasp fundamental principles underlying even simple physics concepts, but it also shows that knowledge injection has reached diminishing marginal returns.
@@
o4-mini o3_o4-mini 46.5 40.5 29.9 35.7 38.2 12.9 35.7 23.2
o3-mini o3-mini 32.9 15.3 16.3 7.1 17.9 53.5 50.4 78.4
o1 o1 38.5 32.0 23.7 23.9 29.5 16.9 38.4 37.9
Gemini-2.5-Pro Gemini-pro 49.0 40.3 32.0 21.0 35.6 17.8 34.7 57.1
Claude-3.7-Sonnet Claude-3.7-Sonnet 27.9 22.8 12.3 20.2 20.8 18.3 55.9 27.6
Doubao-1.5-pro Doubao-1.5-pro 39.0 30.1 24.1 23.9 29.3 22.8 38.2 38.7
GPT-4.1 GPT-4.1 29.2 26.5 18.5 20.4 23.7 9.2 36.6 30.1
GPT-4o GPT-4o 17.1 15.9 10.1 12.7 14.0 7.0 40.9 25.7
QvQ-72B-preview QVQ-72b-preview 16.5 15.3 11.6 15.3 14.7 7.3 29.7 7.3
Qwen2.5-VL-72B Qwen2.5-VL 18.0 16.4 12.1 9.3 13.0 8.9 32.8 48.3
Qwen2.5-VL-7B Qwen2.5-VL 9.6 7.7 5.7 5.2 7.1 19.8 40.6 45.8
Qwen2.5-VL-3B Qwen2.5-VL 6.8 6.8 6.9 3.3 7.1 0.0 -1.5 51.5
@@
o4-mini o3_o4-mini 68.4 68.0 66.7 58.4 65.4 0.6 2.5 14.6
o3-mini o3-mini 62.4 33.6 44.0 12.8 38.2 46.2 29.5 79.5
o1 o1 67.0 64.2 60.8 49.8 60.5 4.2 9.3 25.7
Gemini-2.5-Pro Gemini-pro 72.4 64.4 68.6 39.6 61.3 11.1 5.2 45.3
Claude-3.7-Sonnet Claude-3.7-Sonnet 53.8 47.6 23.6 41.6 41.7 11.5 56.1 22.7
Doubao-1.5-pro Doubao-1.5-pro 68.8 64.8 63.2 41.4 59.6 5.8 8.1 39.8
GPT-4.1 GPT-4.1 53.6 54.0 54.2 28.2 47.5 -0.7 -1.1 47.4
GPT-4o GPT-4o 36.4 40.0 35.4 24.0 34.0 -9.9 2.7 34.1
QvQ-72B-preview QVQ-72b-preview 40.6 38.2 37.1 38.2 38.5 5.9 8.6 5.9
Qwen2.5-VL-72B Qwen2.5-VL 42.8 40.0 38.6 22.0 31.3 6.5 9.8 48.6
Qwen2.5-VL-7B Qwen2.5-VL 17.4 16.2 16.2 9.8 14.9 6.9 6.9 43.7
Qwen2.5-VL-3B Qwen2.5-VL 18.8 15.2 14.2 8.4 14.2 19.1 24.5 55.3

### v8

@@ Abstract
We present SeePhys, a large-scale multimodal benchmark for LLM reasoning grounded in physics questions spanning across middle school to PhD candidacy exams. The benchmark covers 7 fundamental physics domains, incorporating 21 categories of highly heterogeneous diagrams. In contrast to prior works where visual elements mainly serve supplementary purposes, our benchmark features a substantial proportion of vision-essential problems (75%) that mandate visual information extraction for problem-solving. Through extensive evaluation, we observe that even the best-performing reasoning models (e.g., Gemini-2.5-pro and OpenAI-o4-mini) achieve sub-60% accuracy on our benchmark. These results reveal fundamental challenges in current large language models’ visual understanding capabilities, particularly in: (i) establishing meaningful connections between diagram interpretation and physics reasoning, and (ii) overcoming their persistent reliance on textual cues as cognitive shortcuts.
@@ 1 Introduction
While mathematical reasoning has been a cornerstone for evaluating the reasoning capability of large language models (LLMs) and multimodal large language models (MLLMs) [8, 14, 11, 18, 21, 6, 23, 5, 50, 42, 45], the realm of natural science, especially the discipline of physics, remains understudied as a even more diverse testbed for complex scientific reasoning. Physics reasoning inherently binds text explanations to real-world visual contexts, exposing critical gaps in their ability to emulate human-like world modeling in the context of physics problem-solving [10, 24, 51].
@@
Frontier models have begun to demonstrate abstract perception of physical scenarios and logical reasoning capabilities [25, 12, 29, 30, 9]. Due to the intrinsic diversity of physics diagrams and the fact that they inherently reflect the law of physics in our world, developing a comprehensive benchmark for evaluating physics reasoning abilities and cross-modal understanding is crucial for enhancing LLMs.
@@
Early research primarily focused on assessing physical commonsense [4] and basic scientific knowledge [22], which was later gradually extended to competition-level and university-level physics problems [34, 10]. Due to the broad knowledge scope and high annotation difficulty inherent, some studies only considered test samples from limited knowledge levels [52, 13, 48]. Furthermore, other works targeted the evaluation of language models and did not incorporate visual information [33, 10, 46]. Notably, compared to the mathematics domain, there has been limited exploration of models’ perceptual differences in processing physics diagrams, despite their richer topological structures.
@@
Figure 1: Overview of SeePhys. It encompasses 7 core physics domains and 21 diagram types, spanning the full knowledge spectrum from middle school to PhD candidacy exams levels.
@@
With these challenges in mind, we introduce SeePhys, to measure LLMs’ capability to visually understand the law of physics. It is a fully multimodal benchmark for evaluating physics reasoning across middle school to PhD qualifying exam levels. SeePhys comprises 2,000 rigorously validated questions collected from open-source textbooks, exercises, examinations, and competitions. These questions span 7 major fields of both classical and modern physics. To assess the extent to which different models rely on visual information for reasoning, we curate two subsets with different degrees of visual information enrichment and additionally compile supplementary copies of 2,000 purely visual instances where all problem statements in texts are presented in picture form. Through meticulous selection of 21 diagram types by domain experts, each problem challenges frontier MLLMs to integrate domain knowledge with visual understanding of physics diagrams (e.g., Feynman diagrams for particle interactions and Circuit diagrams for Electromagnetism). With SeePhys, we conduct extensive experiments to evaluate 28 leading LLMs and MLLMs such as o4-mini [30] and Gemini-2.5-Pro [9]. The results reveal that even with extensive chain-of-thought, none of the current models could surpass 55% accuracy. Our analysis reveals that even non-essential diagrams can enhance physics reasoning performance when presented to MLLMs. We have also raised a challenge [^1] https://www.codabench.org/competitions/7925/ in the 2nd AI for Math Workshop at ICML 2025222https://sites.google.com/view/ai4mathworkshopicml2025/.
@@ 2 Related Work
Math reasoning benchmarks. Mathematical reasoning has emerged as a central testbed for evaluating LLMs. Early benchmarks like GSM8K [8] established the foundation for multi-step textual reasoning through elementary problems, while MATH [14] advanced the field with competition-level tasks (e.g., AMC/AIME), exposing critical limitations of early models. As these benchmarks achieved saturation, the community shifted toward higher complexity, e.g., introducing Olympiad-level challenges requiring formal theorem proving and combinatorial reasoning [11, 18]. Concurrently, the rise of multimodal reasoning spurred benchmarks such as MathVista [21] and MATH-V [42] to integrate visual comprehension (e.g., diagrams, plots) with compositional reasoning. However, MathVerse [50] has found that MLLMs tend to rely on the reasoning capabilities of language models when performing mathematical tasks. In contrast, scientific diagrams, which are abstractions derived from real-world scenarios, with their complex visual features, may provide a more effective testbed for benchmarking models’ multimodal capabilities.
Physics Reasoning benchmarks. Contemporary physics benchmarks target broad scope of domains, with frontier datasets now encompassing: (1) PhD candidacy exam problems, and (2) Olympiad-style challenges. Text-based physics benchmarks like PHYBench [33] and UGPhysics [46] provide challenging problems that test advanced reasoning skills, yet fundamentally lack the visual components necessary to assess diagram interpretation abilities. Multimodal physics benchmarks such as PhysReason [52], OlympiadBench [13], and PHYSICS [10] emphasize visual reasoning challenges without analysis regarding the extent of visual components’ impact. Moreover, constrained by the high annotation costs stemming from the extensive domain knowledge required and the scarcity of qualified multimodal materials, these datasets lack comprehensive coverage across knowledge hierarchies and detailed annotations of diagram types. To address these limitations, we contribute a dataset with fully multimodal and full-spectrum physics problems.
@@ 3.1 Data Collection Principles
UGPhysics [46] 0 11,040
PHYSICS [10]* 298 1,297
PHYBench [33] 0 500
PhysReason [52] 972 1,200
@@
ScienceQA [22] 2,797 3,215
OlympiadBench [13] 1,958 2,334
SciBench [44]* 177 291
SciEval [37] 0 1,657
MMMU [48] 444 443
MMMU-Pro [49] 65 60
GPQA [34] 0 227
ARB [35] 31 129
HLE [32] 28 230
@@
Wide knowledge spectrum. To provide a comprehensive evaluation of the model’s physics knowledge comprehension and modeling capabilities, questions are systematically selected across the following eight progressive knowledge levels: middle school, high school, beginner Olympiad problems, advanced Olympiad problems, undergraduate, senior undergraduate, master, and PhD candidacy exams. Domain experts in physics are invited to selectively curate the most representative problems at each knowledge tier.
@@ 3.2 Annotation
We then utilize Mathpix [^3] https://mathpix.com/ to perform OCR parsing on the collected documents and convert them into markdown text. Afterwards, we construct the (question, reasoning, answer) triples for each question. GPT-4.1 [27] is employed to process redundant line breaks, string omissions, and LaTeX syntax errors. Finally, the trained annotators use a LaTeX parser to conduct manual verification on the triples.
@@
Data leakage prevention. To minimize the risk of data leakage, we eliminate samples with inconsistent responses by toggling the search function of GPT-4o [28] on and off. Inspired by Rein et al. [34], we subsequently conduct a manual search for the remaining questions with correct responses.
Multimodal enhancement. To further eliminate the influence of textual modality information, we first use o4-mini [30] to add a caption field to each example, containing a detailed description of geometric features and numerical information. We then augment the original 2,000 questions to obtain a purely visual version. Specifically, we render each question along with its corresponding diagrams into a single image up to 4,096×4,096 pixels. During rendering, we introduce variations in font types and sizes while ensuring readability based on the diagram’s dimensions. The introduction of purely visual QA not only enhances the authenticity of evaluation but also advances the model’s cross-modal understanding capabilities—mirroring human cognition by extracting key features from pixels, associating abstract concepts, and ultimately achieving problem-solving accuracy comparable to that under text-image QA conditions. Appendix A demonstrates the cases.
@@ 4.1 Evaluation Protocol
Composite judgment strategy. Recent advancements in the reasoning capabilities of LLMs enable them to discern key information within complex responses, thereby mitigating the inaccuracies often introduced by applying template matching to open-ended questions. Therefore, we further develop a composite judgment strategy based on a combination of LLM and template matching. As a first step, the model generates a response to the given input question and significant figures. Subsequently, the final answer is extracted through template matching and LLM-based processing. During the scoring process, SymPy [^4] https://www.sympy.org is first utilized to perform an initial screening for straightforward answers. Responses that do not pass the screening are then compared with the ground-truth using LLM. We apply accuracy as the metric for this deterministic evaluation. In the experiments in this paper, we use DeepSeek-V3 [16] as the extraction and judge model.
@@ 4.2 Evaluation Models
9 Large language models: DeepSeek-R1 [12], DeepSeek-V3 [16], Qwen3-235B-A22B [40], Qwen2.5-72B-Instruct [47], QwQ-32B [41], R1-Distilled-Llama-70B [12], Llama-4-Scout-17B [26], Gemma3-27B [38], Llama-3.1-8B [20]. We evaluate them with the TC setting.
19 Multimodal large language models: OpenAI o4-mini [30], OpenAI o3-mini [31], OpenAI o1 [29], Gemini-2.5-Pro [9], Claude 3.7 Sonnet [2], Doubao-1.5-pro [36], GPT-4.1 [27], GPT-4o [28], QvQ-72B-preview [39], Qwen-VL series [3, 43], Llama-3.2-Vision series [19], LLaVA-NeXT-7B [17], Phi-4-multimodal [1], InternVL2.5-8B [7], LLaVA-OneVision-7B [15]. All these MLLMs are benchmarked with the TV/TC/TO/VO settings.
@@ 4.3 Performance across Differential Knowledge Levels
DeepSeek-R1 [12] 54.9 46.9 47.7 31.9 49.9 34.2 49.0 41.2 42.2
DeepSeek-V3 [16] 53.9 42.6 36.4 22.8 45.4 29.7 35.9 37.5 36.0
R1-Distilled-Llama-70B [12] 48.0 41.4 34.6 14.2 31.5 16.0 28.9 25.9 26.9
Qwen3-235B-A22B [40] 47.1 33.7 31.8 20.4 41.2 25.1 31.7 30.7 31.1
Qwen2.5-72B-Inst [47] 41.2 40.2 25.2 8.2 26.8 12.8 18.6 17.8 21.1
QwQ-32B [41] 47.1 42.2 44.9 15.5 40.0 20.1 32.4 24.0 29.7
Llama-4-Scout-17B [26] 48.0 36.5 31.8 11.3 28.5 14.2 28.3 26.1 24.8
Gemma3-27B [38] 21.6 36.5 30.8 5.1 23.1 9.1 15.2 11.9 16.9
Llama-3.1-8B [20] 26.5 15.7 17.8 3.9 7.6 3.7 10.3 8.4 9.2
@@
o4-mini [30] 66.7 61.8 56.1 41.8 53.8 45.7 51.0 53.4 51.9
o3-mini [31] 47.1 46.2 39.3 28.3 47.0 36.1 48.3 42.3 40.3
o1 [29] 60.8 56.6 50.5 32.5 54.4 40.6 52.4 40.4 45.6
Gemini-2.5-Pro [9] 69.6 66.7 64.5 46.7 64.2 50.2 53.8 44.2 54.9
Claude-3.7-Sonnet [2] 52.9 51.8 43.0 16.7 41.4 26.5 33.8 32.4 34.6
Doubao-1.5-pro [36] 70.6 58.2 49.5 29.2 56.6 34.7 40.7 37.5 43.9
GPT-4.1 [27] 51.0 52.6 41.1 17.0 39.7 31.1 42.1 35.6 35.3
GPT-4o [28] 37.3 39.0 34.6 7.5 23.4 15.5 24.1 21.8 21.9
QVQ-72b-preview [39] 38.2 36.5 30.8 11.3 25.9 14.2 26.2 20.2 22.5
Qwen2.5-VL-72B-Inst [3] 61.8 42.2 29.0 10.4 29.9 14.6 18.6 19.4 24.2
Qwen2.5-VL-7B-Inst [3] 39.2 25.3 21.5 4.2 8.7 5.9 10.3 7.3 11.6
Qwen2.5-VL-3B-Inst [3] 30.4 21.3 13.1 2.9 10.4 7.3 6.2 6.2 9.8
Qwen2-VL-7B-Inst [43] 24.5 17.3 14.0 4.4 8.5 4.6 10.3 7.0 9.2
Llama-3.2-90B-Vision [19] 21.6 25.7 22.4 3.9 9.3 10.0 12.4 8.9 11.7
Llama3.2-11B-Vision [19] 23.5 18.5 14.0 4.2 5.4 3.7 4.8 7.5 8.3
LLaVA-NeXT-7B [17] 14.5 12.7 11.2 5.5 13.2 8.2 11.0 9.4 8.7
Phi-4-multimodal [1] 20.6 12.4 12.1 4.4 7.0 5.0 8.3 4.9 7.6
InternVL2.5-8B [7] 17.6 12.4 9.3 2.9 5.6 3.2 4.1 5.1 6.2
LLaVA-OneVision-7B [15] 20.6 10.8 12.1 2.7 5.4 2.3 6.2 5.4 6.1
@@
Diminishing returns of knowledge injection. Table 2 further illustrates performance disparities across models at varying knowledge levels. Contrary to expectations, the difficulty progression for models does not strictly follow the knowledge level (e.g., senior undergraduate and advanced Olympiad questions exhibit lower accuracy than PhD candidacy exams). This discrepancy suggests that current models primarily rely on knowledge memorization rather than truly learning the derivation patterns of scientific laws. Furthermore, by examining results from middle school to PhD, we observe that weaker models demonstrate a substantially steeper performance reduction ratio (e.g., LLaVA-OneVision-7B, -74.5%) compared to stronger ones (e.g., o4-mini, -19.9%). Not only indicates that cutting-edge models fail to grasp fundamental principles underlying even simple physics concepts, but it also shows that knowledge injection has reached diminishing marginal returns.
@@
o4-mini [30] 46.5 40.5 29.9 35.7 38.2 12.9 35.7 23.2
o3-mini [31] 32.9 15.3 16.3 7.1 17.9 53.5 50.4 78.4
o1 [29] 38.5 32.0 23.7 23.9 29.5 16.9 38.4 37.9
Gemini-2.5-Pro [9] 49.0 40.3 32.0 21.0 35.6 17.8 34.7 57.1
Claude-3.7-Sonnet [2] 27.9 22.8 12.3 20.2 20.8 18.3 55.9 27.6
Doubao-1.5-pro [36] 39.0 30.1 24.1 23.9 29.3 22.8 38.2 38.7
GPT-4.1 [27] 29.2 26.5 18.5 20.4 23.7 9.2 36.6 30.1
GPT-4o [28] 17.1 15.9 10.1 12.7 14.0 7.0 40.9 25.7
QvQ-72B-preview [39] 16.5 15.3 11.6 15.3 14.7 7.3 29.7 7.3
Qwen2.5-VL-72B [3] 18.0 16.4 12.1 9.3 13.0 8.9 32.8 48.3
Qwen2.5-VL-7B [3] 9.6 7.7 5.7 5.2 7.1 19.8 40.6 45.8
Qwen2.5-VL-3B [3] 6.8 6.8 6.9 3.3 7.1 0.0 -1.5 51.5
@@
o4-mini [30] 68.4 68.0 66.7 58.4 65.4 0.6 2.5 14.6
o3-mini [31] 62.4 33.6 44.0 12.8 38.2 46.2 29.5 79.5
o1 [29] 67.0 64.2 60.8 49.8 60.5 4.2 9.3 25.7
Gemini-2.5-Pro [9] 72.4 64.4 68.6 39.6 61.3 11.1 5.2 45.3
Claude-3.7-Sonnet [2] 53.8 47.6 23.6 41.6 41.7 11.5 56.1 22.7
Doubao-1.5-pro [36] 68.8 64.8 63.2 41.4 59.6 5.8 8.1 39.8
GPT-4.1 [27] 53.6 54.0 54.2 28.2 47.5 -0.7 -1.1 47.4
GPT-4o [28] 36.4 40.0 35.4 24.0 34.0 -9.9 2.7 34.1
QvQ-72B-preview [39] 40.6 38.2 37.1 38.2 38.5 5.9 8.6 5.9
Qwen2.5-VL-72B [3] 42.8 40.0 38.6 22.0 31.3 6.5 9.8 48.6
Qwen2.5-VL-7B [3] 17.4 16.2 16.2 9.8 14.9 6.9 6.9 43.7
Qwen2.5-VL-3B [3] 18.8 15.2 14.2 8.4 14.2 19.1 24.5 55.3